# Holistic Unlearning Benchmark: A Multi-Faceted Evaluation for Text-to-Image Diffusion Model Unlearning

### v2

@@ Abstract
As text-to-image diffusion models gain widespread commercial applications, there are increasing concerns about unethical or harmful use, including the unauthorized generation of copyrighted or sensitive content. Concept unlearning has emerged as a promising solution to these challenges by removing undesired and harmful information from the pre-trained model. However, the previous evaluations primarily focus on whether target concepts are removed while preserving image quality, neglecting the broader impacts such as unintended side effects. In this work, we propose Holistic Unlearning Benchmark (HUB), a comprehensive framework for evaluating unlearning methods across six key dimensions: faithfulness, alignment, pinpoint-ness, multilingual robustness, attack robustness, and efficiency. Our benchmark covers 33 target concepts, including 16,000 prompts per concept, spanning four categories: Celebrity, Style, Intellectual Property, and NSFW. Our investigation reveals that no single method excels across all evaluation criteria. By releasing our evaluation code and dataset, we hope to inspire further research in this area, leading to more reliable and effective unlearning methods.
@@ 1 Introduction
Methods $\mathtt{AC}$ I, S, C, N, O 10
$\mathtt{ESD}$ S, C, N, O 1
$\mathtt{UCE}$ S, C, N, O 1
$\mathtt{SA}$ C, N 50
$\mathtt{RECELER}$ N, O 50
$\mathtt{MACE}$ S, C, N, O 5
Benchmarks I2P ($\mathtt{SLD}$) N 4,703
Ring-A-Bell N 345
CPDM I, S, C 1
UnlearnCanvas S, O 1
@@
Table 1: Comparison with evaluation settings of previous methods and benchmarks. We use the abbreviations I, S, C, N, and O to denote Intellectual Property ($\mathtt{IP}$), artist style ($\mathtt{Style}$), $\mathtt{Celebrity}$, $\mathtt{NSFW}$, and $\mathtt{Object}$, respectively. ✓ indicates that the method quantitatively evaluates the corresponding task. For unlearning methods, we report the prompts per concept as the total number of prompts utilized for I, S, C, and O, excluding N, because all methods use the I2P dataset [38] for $\mathtt{NSFW}$.
Text-to-image diffusion models have achieved remarkable success in various real-world applications, owing to the extensive text and image pairs used during training [32, 7, 28, 30]. However, these pairs are often collected from the Internet, where they may include not only violent, harmful, or unethical materials but also copyrighted or protected intellectual property (IP) [39]. Such datasets give rise to multiple concerns. First, the presence of violent or hateful content can lead to models that produce malicious or ethically problematic outputs [33, 25, 17, 46]. Second, the unauthorized use of copyrighted or trademarked images in training raises legal and ethical issues about ownership and misuse [26]. While many systems deploy safety filters to block unwanted or infringing outputs [35, 34, 32], these filters heavily rely on predefined malicious or protected patterns, making them vulnerable to circumvention-based prompts.
Unlearning methods [9, 10, 20, 13, 8, 16] offer an alternative approach by removing specific target concepts from the model itself. Although these methods are promising, most evaluations have been limited to confirming the absence of target concepts and ensuring acceptable visual quality. This narrow scope often neglects key considerations such as unintended side effects or performance drops on unrelated concepts. Without a more comprehensive framework, it remains difficult to systematically compare unlearning methods and address questions about their effectiveness and limitations.
@@
Using HUB, we evaluate seven recent unlearning methods, including $\mathtt{SLD}$ [38], $\mathtt{AC}$ [20], $\mathtt{ESD}$ [9], $\mathtt{UCE}$ [10], $\mathtt{SA}$ [13], $\mathtt{RECELER}$ [16], and $\mathtt{MACE}$ [24]. Our findings show that no single method outperforms in all perspectives, emphasizing the need for more holistic unlearning approaches. By releasing both the benchmark framework and associated datasets, we aim to illuminate current limitations and inspire new research on effective and reliable unlearning methods.
@@ 2 Related Work
There is a growing body of research on unlearning techniques for pre-trained text-to-image models, aiming to mitigate the generation of specific target concepts. $\mathtt{SLD}$ [38] uses negative prompts to prevent the generation of the target concept. $\mathtt{AC}$ [20] proposes a fine-tuning method that maps the target concept to alternative concepts. $\mathtt{ESD}$ [9] is a fine-tuning method that inversely guides the model against generating a specified target concept text. $\mathtt{UCE}$ [10] updates cross-attention layers with closed-form solutions for unlearning. $\mathtt{SA}$ [13] introduces an unlearning based on continual learning, and $\mathtt{RECELER}$ [16] uses an adapter and a masking scheme. $\mathtt{MACE}$ [24] utilizes masks that identify regions in the input image corresponding to the target concept to guide the unlearning process.
A number of methods have been proposed to evaluate the unlearning methods of text-to-image diffusion models. Several studies focus on optimizing prompts to generate undesired concepts [29, 25, 42, 5, 45, 33, 46]. Additionally, benchmarks have been introduced to assess the effectiveness of unlearning methods [26, 48, 38]. Schramowski et al. [38] propose the I2P dataset to assess the ability of a model to avoid generating inappropriate content that could be offensive, insulting, or anxiety-inducing. Zhang et al. [48] introduce a stylized image dataset to evaluate models that have undergone style unlearning. Similarly, Ma et al. [26] offer a copyright dataset to measure how effectively an unlearned model protects copyrighted material by not reproducing protected content.
@@ 3.1 Concepts Categorization and Detection
Attack robustness We evaluate whether the unlearning method is resistant to adversarial prompts that attempt to recover the forgotten concepts. This includes testing the model with paraphrased, indirect, or obfuscated queries that might bypass unlearning constraints. 1
@@
Concept detection. Through the benchmark, multiple tasks require detecting the presence of a specific concept given a generated image. To identify the presence of the concepts in images, we employ specialized classifiers if a target concept classifier exists. For example, for $\mathtt{NSFW}$ identification, we use Q16 [37], a CLIP-based classifier specifically designed to detect inappropriate content. For identifying celebrities, we use the GIPHY celebrity detector [12], known to detect the faces of celebrities with 98% accuracy. The classifier also learns the faces of the ten celebrities used in this study.
For the concepts of which pretrained classifiers are unavailable, we propose a concept detection framework based on vision-language models (VLMs). Specifically, the detection process consists of two steps combining in-context learning [3] and chain of thought [43]. In the first step, a VLM is provided with three images generated from a reference model with a prompt describing a concept, allowing the VLM to recognize the target concept through in-context learning. In the second step, a test image, generated from a non-reference model with the same prompt used to generate the reference image, is provided and analyzed through the chain of thought reasoning to determine whether the target concept is present. In many of the following scenarios, the reference model is the baseline model before unlearning, and the non-reference models are the unlearned models from the baseline. This approach enables flexible and concept-agnostic concept detection without the need for dedicated classifiers.
We evaluate our detection framework using InternVL [4] and Qwen [1] as backbone VLMs. Specifically, we test the detection framework on two datasets: the Disney character dataset [36] and AI-ArtBench [41] corresponding to $\mathtt{IP}$ and $\mathtt{Style}$, respectively. We measure performance by calculating the percentage of correctly identified images. On average, our detection framework with InternVL2.5-8B achieves an accuracy of 83.2% on $\mathtt{IP}$ and 82.5% on $\mathtt{Style}$. With Qwen2.5-VL-7B, the accuracies are 85.1% and 76.1% for $\mathtt{IP}$ and $\mathtt{Style}$, respectively. For the rest of the paper, we report detection performance using InternVL as the backbone VLM unless otherwise noted. Additional details, including experimental settings, can be found in Appendix B.
@@ 3.2.1 Faithfulness
General image quality. The unlearning process may influence the generation process of the target unrelated concepts. To access the image quality of the target unrelated concepts, we measure the Fréchet Inception Distance (FID) [14] between real COCO images and images generated from unlearned models. We also compute the FID between the original and unlearned models, which we call FID-SD [10]. We use 30k captions from the MS-COCO [22] dataset as the target-unrelated prompts.
Target image quality. Although the general image quality metric gives a broad view of how well the model generates images overall, it may not capture quality losses specific to images generated by prompts containing target concepts. Unlike the general quality, a statistical metric such as FID cannot be used to measure the quality of generated images with target concepts. To address this issue, we evaluate the quality of images generated with the target concept prompts using an aesthetic score proposed in Schuhmann et al. [39], which measures the quality of an individual image. Aesthetic score allows us to isolate and measure any visual degradation that might arise in images explicitly tied to the target concept.
@@ 3.2.2 Alignment
General alignment. The alignment between text prompts and generated images is widely studied in previous works such as Kirstain et al. [19] and Xu et al. [44]. The general alignment task measures how unlearning influences the overall alignments between prompts and images. To do this, we generate images using 30k captions from the MS-COCO dataset and measure the alignment scores via PickScore [19] and ImageReward [44].
@@
Recent research uses the QG/A (question generation and answering) framework to evaluate text-to-image model alignment [15, 6, 47]. Building on this approach, we design a QG/A framework to quantitatively evaluate the selective alignment performance. Suppose that we are given a prompt that contains multiple entities, including the target concept. Using an LLM, we first extract all explicitly mentioned physical entities from the prompt, excluding the target concept. Subsequently, we formulate a question for each entity to verify its presence in the image. We pass these questions into a VLM with the generated image to measure the proportion of affirmative responses. We randomly select 1,000 prompts for each target concept from the set generated in Sec. 3.1. We do not conduct experiments specifically on $\mathtt{NSFW}$ content, as $\mathtt{NSFW}$ concepts tend to influence the overall prompt, confounding our results.
@@ 3.2.3 Pinpoint-ness
Unlearning a specific concept may unintentionally affect similar but non-target concepts. For example, removing Mickey Mouse might lead the model to forget Minnie Mouse, causing an over-erasing effect. The pinpoint-ness perspective evaluates how accurately an unlearning method removes the target concept while minimizing unintended effects. $\mathtt{MACE}$ [24] also addresses a similar problem but evaluates only predefined lexicons (e.g., for $\mathtt{Style}$, it considers other artist styles). In contrast, we leverage the shared feature representations inherent in CLIP models [31].
We pick 100 lexicons from WordNet [27] with the highest CLIP scores for each target concept. Then, we generate 10 images for each lexicon with a simple but straightforward prompt “a photo of {lexicon}” and report the proportion of images containing the target lexicon.
@@ 3.2.4 Multilingual robustness
$\mathtt{Celebrity}$ $\mathtt{Original}$ 0.628 13.203 5.433 0.172 0.555 0.482 0.686 0.437 0.0
@@
$\mathtt{Style}$ $\mathtt{Original}$ 0.638 13.203 5.586 0.172 0.570 0.698 0.438 0.339 0.0
@@
$\mathtt{IP}$ $\mathtt{Original}$ 0.683 13.203 5.253 0.172 0.566 0.578 0.510 0.393 0.0
$\mathtt{SLD}$ 0.349 15.932 5.189 0.089 0.573 0.525 0.150 0.164 0.0
@@
$\mathtt{MACE}$ 0.050 12.988 5.229 -0.007 0.594 0.383 0.057 0.033 137.1
$\mathtt{NSFW}$ $\mathtt{Original}$ 0.647 13.203 5.100 0.172 0.609 0.322 0.796 0.0
$\mathtt{SLD}$ 0.339 17.838 5.319 0.107 0.541 0.080 0.506 0.0
$\mathtt{AC}$ 0.438 16.394 4.955 0.116 0.453 0.178 0.588 59.6
$\mathtt{ESD}$ 0.343 15.733 5.115 -0.284 0.124 0.159 0.476 106.0
$\mathtt{UCE}$ 0.603 13.954 5.076 0.190 0.571 0.248 0.780 0.1
$\mathtt{SA}$ 0.327 53.384 4.839 -0.781 0.097 0.126 0.447 34165.0
$\mathtt{RECELER}$ 0.272 15.882 5.192 -0.066 0.327 0.085 0.389 100.0
$\mathtt{MACE}$ 0.344 22.153 4.856 -1.403 0.133 0.354 0.360 150.7
Overall $\mathtt{Original}$ 0.649 13.203 5.343 0.172 0.564 0.592 0.489 0.491 0.0
$\mathtt{SLD}$ 0.228 16.567 5.296 0.077 0.569 0.502 0.086 0.196 0.0
$\mathtt{AC}$ 0.301 14.203 5.308 0.127 0.601 0.528 0.198 0.280 37.3
$\mathtt{ESD}$ 0.143 14.525 5.265 -0.082 0.548 0.260 0.069 0.148 106.0
$\mathtt{UCE}$ 0.250 13.822 5.339 0.193 0.577 0.535 0.116 0.252 0.1
$\mathtt{SA}$ 0.173 32.572 5.102 -0.322 0.508 0.131 0.080 0.166 29980.0
$\mathtt{RECELER}$ 0.086 15.190 5.270 0.006 0.497 0.316 0.028 0.107 100.0
$\mathtt{MACE}$ 0.148 15.303 5.262 -0.345 0.566 0.306 0.122 0.125 140.3

### v3

@@ Abstract
As text-to-image diffusion models gain widespread commercial applications, there are increasing concerns about unethical or harmful use, including the unauthorized generation of copyrighted or sensitive content. Concept unlearning has emerged as a promising solution to these challenges by removing undesired and harmful information from the pre-trained model. However, the previous evaluations primarily focus on whether target concepts are removed while preserving image quality, neglecting the broader impacts such as unintended side effects. In this work, we propose Holistic Unlearning Benchmark (HUB), a comprehensive framework for evaluating unlearning methods across six key dimensions: faithfulness, alignment, pinpoint-ness, multilingual robustness, attack robustness, and efficiency. Our benchmark covers 33 target concepts, including 16,000 prompts per concept, spanning four categories: Celebrity, Style, Intellectual Property, and NSFW. Our investigation reveals that no single method excels across all evaluation criteria. By releasing our evaluation code and dataset [^1] https://github.com/ml-postech/HUB, we hope to inspire further research in this area, leading to more reliable and effective unlearning methods.
@@ 1 Introduction
Methods
$\mathtt{AC}$ [kumari2023ablating] I, S, C, N, O 10
$\mathtt{ESD}$ [gandikota2023erasing] S, C, N, O 1
$\mathtt{UCE}$ [gandikota2024unified] S, C, N, O 1
$\mathtt{SA}$ [heng2024selective] C, N 50
$\mathtt{RECELER}$ [huang2023receler] N, O 50
$\mathtt{MACE}$ [lu2024mace] S, C, N, O 5
Benchmarks
I2P ($\mathtt{SLD}$[schramowski2023safe] N 4,703
Ring-A-Bell [ringabell] N 345
CPDM [ma2024dataset] I, S, C 1
UnlearnCanvas [zhang2024unlearncanvas] S, O 1
@@ Methods
Table 1: Comparison with evaluation settings of previous methods and benchmarks. We use the abbreviations I, S, C, N, and O to denote Intellectual Property ($\mathtt{IP}$), artist style ($\mathtt{Style}$), $\mathtt{Celebrity}$, $\mathtt{NSFW}$, and $\mathtt{Object}$, respectively. ✓ indicates that the method quantitatively evaluates the corresponding task. For unlearning methods, we report the prompts per concept as the total number of prompts utilized for I, S, C, and O, excluding N, because all methods use the I2P dataset [schramowski2023safe] for $\mathtt{NSFW}$.
Text-to-image diffusion models have achieved remarkable success in various real-world applications, owing to the extensive text and image pairs used during training [ramesh2022hierarchical, dhariwal2021diffusion, nichol2021glide, podell2023sdxl]. However, these pairs are often collected from the Internet, where they may include not only violent, harmful, or unethical materials but also copyrighted or protected intellectual property (IP) [schuhmann2022laion]. Such datasets give rise to multiple concerns. First, the presence of violent or hateful content can lead to models that produce malicious or ethically problematic outputs [rando2022red, ma2024jailbreaking, kim2024automatic, yang2024sneakyprompt]. Second, the unauthorized use of copyrighted or trademarked images in training raises legal and ethical issues about ownership and misuse [ma2024dataset]. While many systems deploy safety filters to block unwanted or infringing outputs [saharia2022photorealistic, rombach2022high, ramesh2022hierarchical], these filters heavily rely on predefined malicious or protected patterns, making them vulnerable to circumvention-based prompts.
Unlearning methods [gandikota2023erasing, gandikota2024unified, kumari2023ablating, heng2024selective, fan2023salun, huang2023receler] offer an alternative approach by removing specific target concepts from the model itself. Although these methods are promising, most evaluations have been limited to confirming the absence of target concepts and ensuring acceptable visual quality. This narrow scope often neglects key considerations such as unintended side effects or performance drops on unrelated concepts. Without a comprehensive framework, it remains difficult to systematically compare unlearning methods and address questions about their effectiveness and limitations.
@@
Using HUB, we evaluate seven recent unlearning methods, including $\mathtt{SLD}$ [schramowski2023safe], $\mathtt{AC}$ [kumari2023ablating], $\mathtt{ESD}$ [gandikota2023erasing], $\mathtt{UCE}$ [gandikota2024unified], $\mathtt{SA}$ [heng2024selective], $\mathtt{RECELER}$ [huang2023receler], and $\mathtt{MACE}$ [lu2024mace]. Our findings show that no single method outperforms in all perspectives, emphasizing the need for more holistic unlearning approaches. By releasing the benchmark framework and associated datasets, we aim to illuminate current limitations and inspire new research on effective and reliable unlearning methods.
@@ 2 Related Work
There is a growing body of research on unlearning techniques for pre-trained text-to-image models, aiming to mitigate the generation of specific target concepts. $\mathtt{SLD}$ [schramowski2023safe] uses negative prompts to prevent the generation of the target concept. $\mathtt{AC}$ [kumari2023ablating] proposes a fine-tuning method that maps the target concept to alternative concepts. $\mathtt{ESD}$ [gandikota2023erasing] is a fine-tuning method that inversely guides the model against generating a specified target concept text. $\mathtt{UCE}$ [gandikota2024unified] updates cross-attention layers with closed-form solutions for unlearning. $\mathtt{SA}$ [heng2024selective] introduces an unlearning based on continual learning, and $\mathtt{RECELER}$ [huang2023receler] uses an adapter and a masking scheme. $\mathtt{MACE}$ [lu2024mace] utilizes masks that identify regions in the input image corresponding to the target concept to guide the unlearning process.
A number of methods have been proposed to evaluate the unlearning methods of text-to-image diffusion models. Several studies focus on optimizing prompts to generate undesired concepts [pham2023circumventing, ma2024jailbreaking, ringabell, chin2023prompting4debugging, yang2024mma, rando2022red, yang2024sneakyprompt]. Additionally, benchmarks have been introduced to assess the effectiveness of unlearning methods [ma2024dataset, zhang2024unlearncanvas, schramowski2023safe]. schramowski2023safe propose the I2P dataset to assess the ability of a model to avoid generating inappropriate content that could be offensive, insulting, or anxiety-inducing. zhang2024unlearncanvas introduce a stylized image dataset to evaluate models that have undergone style unlearning. Similarly, ma2024dataset offer a copyright dataset to measure how effectively an unlearned model protects copyrighted material by not reproducing protected content.
@@ 3.1 Concepts Categorization and Detection
Attack robustness We evaluate whether the unlearning method is resistant to adversarial prompts that attempt to recover the forgotten concepts. This includes testing the model with paraphrased, indirect, or obfuscated queries that might bypass unlearning constraints. 3
@@
Concept detection. Through the benchmark, multiple tasks require detecting the presence of a specific concept given a generated image. To identify the presence of the concepts in images, we employ specialized classifiers if a target concept classifier exists. For example, for $\mathtt{NSFW}$ identification, we use Q16 [schramowski2022can], a CLIP-based classifier specifically designed to detect inappropriate content. For identifying celebrities, we use the GIPHY celebrity detector [GCD], known to detect the faces of celebrities with 98% accuracy. The classifier also learns the faces of the ten celebrities used in this study.
For the concepts of which pretrained classifiers are unavailable, we propose a concept detection framework based on vision-language models (VLMs). Specifically, the detection process consists of two steps combining in-context learning [brown2020language] and chain of thought [wei2022chain]. In the first step, a VLM is provided with three images generated from a reference model with a prompt describing a concept, allowing the VLM to recognize the target concept through in-context learning. In the second step, a test image, generated from a non-reference model with the same prompt used to generate the reference image, is analyzed via the chain of thought reasoning to determine whether the target concept is present. In many of the following scenarios, the reference model is the baseline model before unlearning, and the non-reference models are the unlearned models from the baseline. This approach enables flexible and concept-agnostic concept detection without the need for dedicated classifiers.
We evaluate our detection framework using InternVL [chen2024expanding] and Qwen [bai2025qwen2] as backbone VLMs. Specifically, we test the detection framework on two datasets: the Disney character dataset [disney_characters_dataset] and AI-ArtBench [ai_artbench] corresponding to $\mathtt{IP}$ and $\mathtt{Style}$, respectively. We measure performance by calculating the percentage of correctly identified images. On average, our detection framework with InternVL2.5-8B achieves an accuracy of 83.2% on $\mathtt{IP}$ and 82.5% on $\mathtt{Style}$. With Qwen2.5-VL-7B, the accuracies are 85.1% and 76.1% for $\mathtt{IP}$ and $\mathtt{Style}$, respectively. For the rest of the paper, we report detection performance using InternVL as the backbone VLM unless otherwise noted. Additional details, including experimental settings, can be found in Appendix B.
@@ 3.2.1 Faithfulness
General image quality. The unlearning process may influence the generation process of the target unrelated concepts. To access the image quality of the target unrelated concepts, we measure the Fréchet Inception Distance (FID) [heusel2017gans] between real COCO images and images generated from unlearned models. We also compute the FID between the original and unlearned models, which we call FID-SD [gandikota2024unified]. We use 30k captions from the MS-COCO [cocodataset] dataset as the target-unrelated prompts.
Target image quality. Although the general image quality metric gives a broad view of how well the model generates images overall, it may not capture quality losses specific to images prompted with target concepts. Unlike the general quality, a statistical metric such as FID cannot measure the quality of generated images with target concepts. To address this issue, we evaluate the quality of images generated with the target concept prompts using an aesthetic score proposed in schuhmann2022laion, which measures the quality of an individual image. Aesthetic score allows us to isolate and measure any visual degradation that might arise in images explicitly tied to the target concept.
@@ 3.2.2 Alignment
General alignment. The alignment between text prompts and generated images is widely studied in previous works such as kirstain2023pick and xu2024imagereward. The general alignment task measures how unlearning influences the overall alignments between prompts and images. To do this, we generate images using 30k captions from the MS-COCO dataset and measure the alignment scores via PickScore [kirstain2023pick] and ImageReward [xu2024imagereward].
@@
Recent research uses the QG/A (question generation and answering) framework to evaluate text-to-image model alignment [hu2023tifa, JaeminCho2024, yarom2023you]. Building on this approach, we design a QG/A framework to quantitatively evaluate the selective alignment performance. Suppose that we are given a prompt that contains multiple entities, including the target concept. Using an LLM, we first extract all explicitly mentioned physical entities from the prompt, excluding the target concept. Subsequently, we formulate a question for each entity to verify its presence in the image. We pass these questions into a VLM with the generated image to measure the proportion of affirmative responses. We randomly select 1,000 prompts for each target concept from the set generated in Sec. 3.1. We do not conduct experiments specifically on $\mathtt{NSFW}$ content, as $\mathtt{NSFW}$ concepts tend to influence the overall prompt, confounding our results.
@@ 3.2.3 Pinpoint-ness
Unlearning a specific concept may unintentionally affect similar but non-target concepts. For example, removing Mickey Mouse might lead the model to forget Minnie Mouse, causing an over-erasing effect. The pinpoint-ness perspective evaluates how accurately an unlearning method removes the target concept while minimizing unintended effects. $\mathtt{MACE}$ [lu2024mace] also addresses a similar problem but evaluates only predefined lexicons (e.g., for $\mathtt{Style}$, it considers other artist styles). In contrast, we leverage the shared feature representations inherent in CLIP models [radford2021learning].
We pick 100 lexicons from WordNet [miller1995wordnet] with the highest CLIP scores for each target concept. Then, we generate 10 images for each lexicon with a simple but straightforward prompt “a photo of {lexicon}” and report the proportion of images containing the target lexicon.
@@ 3.2.4 Multilingual robustness
$\mathtt{Celebrity}$
$\mathtt{Original}$ 0.628 13.203 5.433 0.172 0.555 0.482 0.686 0.437 0.0
@@
$\mathtt{Style}$
$\mathtt{Original}$ 0.638 13.203 5.586 0.172 0.570 0.698 0.438 0.339 0.0
@@
$\mathtt{IP}$
$\mathtt{Original}$ 0.683 13.203 5.253 0.172 0.566 0.578 0.510 0.393 0.0
$\mathtt{SLD}$ 0.349 15.932 5.189 0.089 0.573 0.525 0.207 0.164 0.0
@@
$\mathtt{MACE}$ 0.050 12.988 5.229 -0.007 0.594 0.383 0.031 0.033 137.1
$\mathtt{NSFW}$
$\mathtt{Original}$ 0.647 13.203 5.100 0.172 0.609 0.322 0.796 0.0
$\mathtt{SLD}$ 0.339 17.838 5.319 0.107 0.541 0.085 0.506 0.0
$\mathtt{AC}$ 0.438 16.394 4.955 0.116 0.453 0.182 0.588 59.6
$\mathtt{ESD}$ 0.343 15.733 5.115 -0.284 0.124 0.167 0.476 106.0
$\mathtt{UCE}$ 0.603 13.954 5.076 0.190 0.571 0.241 0.780 0.1
$\mathtt{SA}$ 0.327 53.384 4.839 -0.781 0.097 0.121 0.447 34165.0
$\mathtt{RECELER}$ 0.272 15.882 5.192 -0.066 0.327 0.093 0.389 100.0
$\mathtt{MACE}$ 0.344 22.153 4.856 -1.403 0.133 0.337 0.360 150.7
Overall
$\mathtt{Original}$ 0.649 13.203 5.343 0.172 0.564 0.592 0.489 0.491 0.0
$\mathtt{SLD}$ 0.228 16.567 5.296 0.077 0.569 0.502 0.101 0.196 0.0
$\mathtt{AC}$ 0.301 14.203 5.308 0.127 0.601 0.528 0.199 0.280 37.3
$\mathtt{ESD}$ 0.143 14.525 5.265 -0.082 0.548 0.260 0.071 0.148 106.0
$\mathtt{UCE}$ 0.250 13.822 5.339 0.193 0.577 0.535 0.114 0.252 0.1
$\mathtt{SA}$ 0.173 32.572 5.102 -0.322 0.508 0.131 0.079 0.166 29980.0
$\mathtt{RECELER}$ 0.086 15.190 5.270 0.006 0.497 0.316 0.030 0.107 100.0
$\mathtt{MACE}$ 0.148 15.303 5.262 -0.345 0.566 0.306 0.111 0.125 140.3